At a certain hill station pure water boils at 99.725 degree celsius. If KB for water is. 513 degree Celsius kg per mole the boiling point of. 69 and solution of Urea will be

Grade:12

1 Answers

Chanchal
5 years ago
Delta tb = kb * m =0.513*0.69= 0.3539degree clecius Bpt. Of solution = 99.725degree celcius+0.3539degree celcius = 100.079degree celcius